A plane flies from base camp to lake A, a distance of 280 km at a direction of 20 degrees north of east. After dropping off supplies, the plane flies to lake B, which is 190km and 30 degree west of north from lake A.





1
Expert's answer
2021-12-16T17:46:09-0500







Analyzing the diagram

Let "x=" the distance between Base Camp and Lake B


Angle "\beta°=180-(70+30)"

"=80\u00b0"


Using cosine rule

"x^2=190^2+280^2-2×190×180\>Cos\>80"

"x=309.88km"



Using Sine Rule

"\frac{309.88}{Sin80}=\frac{280}{Sin({\theta}+30)}"


Sin "(\theta+30)=0.8898"

"\theta=32.85"


To return to Base the plan flies "309.88km" at a direction of "32.85\u00b0" west of south from Lake B